Consider Peter’s Pizza Place. Peter needs to know if he is optimizing cost in handling his inputs of capital and labour. This is what he sees: (5)

Marginal Product of Capital 3,200 pizzas/oven
Marginal Product of Labour 1,400 pizzas/worker
Wage Rate $280/week
Rental Price of Ovens $800/week

Using this information, find out if Peter is optimizing his costs and if he is not, then how should he reallocate his factors of production?
1
Expert's answer
2018-12-25T09:53:11-0500

Cost optimization condition is:

MPL/w = MPk/r,

1400/280 = 3200/800,

5 = 4, so Peter is not optimizing his costs and he should reallocate his factors of production and use more units of capital and less units of labour.